Shift Summary: 10/18/99 Owl

Jason LaDue (jladue@cuhep.Colorado.EDU)
Mon, 18 Oct 1999 15:54:53 -0600 (MDT)

Shift Summary for owl shift, 10/16/99

On Shift: Bob Hsiung, J. LaDue
Also Appearing: Bob T., Sydney

Hours of beam available at all intensities: 3.45
Hours of beam available at nominal intensity: 3.45
Hours of E799 data (special runs not included): 2.75
Hours of special runs (laser scans, muon runs...): 0.75

Accounting of lost time:
Stopping/starting runs.

Runs ended this shift:
14861 - 3.6M - BEAM_799
14862 - 4.7M - BEAM_799
14863 - - LASER SCAN

Tapes brought to FCC this shift:

Accesses during shift:
Interlocks dropped because of long down time.

Overview of shift:
Beam goes away because of scheduled downtime. We start using
the new l3 processor scheme with apparent success. Everything is turn-
ed off because of power outage.

Description of Shift:

run spill time comment
14681 00:23 Run started with one less proccess on the filter
planes. This causes a net increase in the pre-
formance of filtering as now gateway has a CPU to
itself. See logbook p 114-5 for explaination.

14681 20 DAQDEAD from ktev3. This is thought to be do to
ktev3 filter code now being forced to run on
individual CPUs.

14681 01:45 Run stopped so that recent changes to ktev3 are
undone. Filter planes are still running one
less proccess.

14682 02:10 Run started.

14682 04:03 Run stopped, beam is gone. Sydney falls asleep
on the couch.

14683 04:15 Laser Scan started, immediately hear from Sydney.

14683 04:55 Laser Scan completes but does not end gracefully.

05:30 Bob T. and Jason L. begin shutting things down.

05:55 All connections to outside world are gone.

06:00 X-Terminals will not reboot.

Follow up Notes on Previous Shift

Items to Follow up on Next Shift